geerky42 (geerky42):

For what value of p does \(\large\displaystyle\sum_{n=1}^{\infty} \dfrac{(-1)^n}{n^p}\) converge?

geerky42 (geerky42):

My guess would be \(0<p<\infty\)

geerky42 (geerky42):

Am I correct?

OpenStudy (agent0smith):

Isn't that just the p-test? p-Series Convergence: http://www.math.com/tables/expansion/tests.htm if p>1 it converges.

geerky42 (geerky42):

Well, with \((-1)^n\), it's different story. Using Alternating series test, \(\sum (-1)^na_n\) converges if \(\lim_{n \rightarrow \infty}a_n = 0\)

OpenStudy (agent0smith):

Yeah i was thinking use the alternating series test... too tired though, i should be asleep :P

geerky42 (geerky42):

\(a_n = \dfrac{1}{n^p}\) so I'm sure it's 0 < p < ∞

geerky42 (geerky42):

Oh lol ok.

OpenStudy (agent0smith):

Yeah it should converge for 0 < p < inf then, since the terms are decreasing.

geerky42 (geerky42):

yeah

geerky42 (geerky42):

Just want to make sure, sorry. Thanks.